Shey Dental Products has current Sales of $540,000, $380,000 of Fixed Assets, and is operating at 85 percent of fixed asset capacity How much Fixed Assets will be necessary if sales grow to $696,600? Round to the nearest whole number and no commas. Answer:
Answer #1

Answer:

Fixed assets for sales of $540000 F=380000*85%=$323000

So Fixed assets for sales of $696600= (696600/540000)*F

Fixed assets for sales of $696600= (696600/540000)*323000=$416,670